1. Albert conducts two studies on the effectiveness of a safe-sex education program. Self-evaluation ratings among participants before, during, and after the program were measured in both studies. In study 1, 13 participants were observed, and in Study 2, 15 participants were observed. If Fobt = 3.38 in both studies, then in which study will the decision be to reject the null hypothesis at a .05 level of significance?

A. Study1   B. Study 2 C.Both D. none

2. A group of 100 Californians completed a questionnaire that measured their belief that the world is fair. A group of 100 New Yorkers completed that same questionnaire. What statistics should you use to determine if there is a difference in the world belief between Californians and New Yorkers?

A. repeated- measures ANOVA B. related sample t C. independent-sample t D. one-way ANOVA E. correlation
Answer #1

1. We have k= 3 groups (ratings of people before, after and during the program)

So k-1=2

For group 1, N= 13+13+13 = 39

So, N-k = 36

So p-value of this ANOVA test is: P(F2,36>3.38) = 0.0452 (from f tables)

Since this is less than 0.05, we have sufficient evidence to reject H0 here

For group 2: k-1=2

N= 15+15+15=45

So N-k = 42

So p-value of this ANOVA test is: P(F2,42>3.38) = 0.0435

Again this is less than the significance level of 0.05, so we have sufficient evidence to reject H0.

So in both the studies we will reject H0.

So option C is correct.

2. We have two independent groups here (100 Californians and 100 new Yorkers). So to test if there's a difference in their opinions we will be using the independent samples t test. So option C is correct.